z1=3.45∠980°

z2=5+9i


  1. z1+z2 in polar form
  2. z1-z2 trigonometric form
  3. z2*z1 exponential form
1
Expert's answer
2022-03-04T06:54:31-0500


z1=3.45∠980° = 3.45 ∠ 260o


= 3.45(cos 260o + j sin 260o)


= −0.6 − 3.4i

z2=5+9i

="Now, \displaystyle{r}=\sqrt{{{\left({5}\right)}^{2}+{\left({9}\right)}^{2}}}=\sqrt{{{106}}}"

"And, \displaystyle\theta= \arctan{{\left(\frac{9}{{5}}\right)}}" = 610

z2 = "\sqrt{{{106}}}" (cos(610)+jsin(610))


(1) z1+z2 in polar form


z1+z2 = (−0.6 − 3.4j) + (5+9j) = 4.4 + 5.6j


In polar form,



"Now, \displaystyle{r}=\sqrt{{{\left({4.4}\right)}^{2}+{\left({5.6}\right)}^{2}}}={7.12}"


"And, \displaystyle\theta= \arctan{{\left(\frac{5.60}{{4.4}}\right)}}"
  • =51.84o


Hence, 4.40 + 5.60j = 7.12 ∠ 51.84o [Answer]


(2) z1-z2 trigonometric form


z1- z2 =( −0.6 − 3.4i) - (5+9i)

= -5.6 - 12.4j


In trigonometric form


"Now, \displaystyle{r}=\sqrt{{{\left({-5.6}\right)}^{2}+{\left({-12.4}\right)}^{2}}}={13.606}"

"And, \displaystyle\theta= \arctan{{\left(\frac{-12.4}{{-5.6}}\right)}}"

= 2460


Thus,

z1- z2 = 13.606 (cos (2460)+jsin(2460)) [Answer]



(3) z2*z1 exponential form


z1=3.45∠980° = "3.45e^{j(\frac{980}{180}\pi )}" = "3.45e^{j(5.44\pi )}"

z2 = 5 +9j = "\sqrt{{{106}}}e^{j(\frac{61}{180}\pi)} = \sqrt{{{106}}}e^{j(0.34\pi)}"


hence,


z2*z1 = "3.45e^{j(5.44\pi )}*\sqrt{{{106}}}e^{j(0.34\pi)}"

="3.45*\sqrt{{{106}}}e^{j(5.44\pi +0.34\pi) )}"


= "35.52e^{j(5.78\pi )}" [Answer]